Manipulale this forau las Er-I 2 to S-Exi-n गो-न

how to you manipulate the first equation to became the second equation
Answer #1

S = E (Xi -3- Now, take 8 X: ) - * - X1 + 2 xix) * * * * * * - - 12 - 19 3- 3 = ซา = 52 x| X 4 ต - x x = - Ex. + ทรา 21 22